APPLICATION OF A DIRECT POTENTIAL FITTING METHOD TO THE B^{1}Σ^{+} AND X^{1}Σ^{+} ELEC-TRONIC STATES OF HF AND DF

JOHN A. COXON, Department of Chemistry, Dalhousie University, Halifax, Nova Scotia B3H 4J3, Canada; \underline{PHOTOS G. HAJIGEORGIOU}, Department of Biosciences, Intercollege, Nicosia, Cyprus.

A collection of 6070 spectroscopic line positions for hydrogen fluoride and deuterium fluoride, which consists of all microwave and infrared X^{1}Σ^{+} ground electronic state data, and the B^{1}Σ^{+} - X^{1}Σ^{+} emission band system data, has been employed in a weighted least-squares fit directly to the radial Hamiltonian operators of the B^{1}Σ^{+} and X^{1}Σ^{+} electronic states. The radial Hamiltonian operator model includes first- and second-order corrections to the Born-Oppenheimer approximation and was derived from the landmark theoretical work of Watson^{a}. The principle isotopomer fitting strategy of Le Roy^{b} was incorporated in the least-squares fit. A total of 54 adjustable parameters was required to obtain a satisfactory representation of the experimental data, with a reduced standard deviation of 1.03. In addition, a collection of highly accurate quantum-mechanically meaningful rotational and centrifugal distortion constants was calculated from the derived Hamiltonian operators of the two electronic states.

aJ. K. G. Watson, J. Mol. Spectrosc. 80, 411 (1980).

bR. J. Le Roy, J. Mol. Spectrosc. 194, 189 (1999).